# Diverse Intra- and Inter-Domain Activity Style Fusion for Cross-Person Generalization in Activity Recognition

### v1

@@ Abstract.
Existing domain generalization (DG) methods for cross-person generalization tasks often face challenges in capturing intra- and inter-domain style diversity, resulting in domain gaps with the target domain. In this study, we explore a novel perspective to tackle this problem, a process conceptualized as domain padding. This proposal aims to enrich the domain diversity by synthesizing intra- and inter-domain style data while maintaining robustness to class labels. We instantiate this concept using a conditional diffusion model and introduce a style-fused sampling strategy to enhance data generation diversity. In contrast to traditional condition-guided sampling, our style-fused sampling strategy allows for the flexible use of one or more random styles to guide data synthesis. This feature presents a notable advancement: it allows for the maximum utilization of possible permutations and combinations among existing styles to generate a broad spectrum of new style instances. Empirical evaluations on a board of datasets demonstrate that our generated data achieves remarkable diversity within the domain space. Both intra- and inter-domain generated data have proven to be significant and valuable, contributing to varying degrees of performance enhancements. Notably, our approach outperforms state-of-the-art DG methods in all human activity recognition tasks.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/XXXXXX.XXXXXXccs: Human-centered computing Ubiquitous computingccs: Computing methodologies Transfer learning
@@
Figure 1. T-SNE visualization (Van der Maaten and Hinton, 2008) of time-series features extracted by various methods across three domains in HAR. Existing representation learning methods result in domain gaps as in both (a) and (b), covering a small portion of target domain (red circles). Standard data augmentation (DA) leads to augmented data (stars), with source domains (orange/blue circles) remaining in close proximity to each other and failing to fill gaps. Our method (d) creates a comprehensive feature space by padding domain gaps via the idea of (e).
@@ 2. Related Work
Human activity recognition (HAR) uses sensors like ECGs for recognizing activities in healthcare and human-computer interaction (Chen et al., 2021; Xu et al., 2023). The complexity of daily activities, varying among individuals with different personal styles, makes recognition challenging (Qian et al., 2021; Zhang et al., 2022a). With the rise of deep learning, deep neural networks have been increasingly utilized to extract informative features from activity signals (Yang et al., 2015; Jiang and Yin, 2015; Zhang et al., 2022a, 2023a). For example, DeepConvLSTM incorporates convolutional and LSTM units for multimodal wearable sensors (Qiao et al., 2020). MultitaskLSTM extracts features using shared weights, then classifies activities and estimates intensity separately (Barut et al., 2020).
@@ 3.1. Problem Statement
In cross-person activity recognition (Qin et al., 2023), a domain is characterized by a joint probability distribution $P_{X,Y}$ across the product space of time-series instances $\mathcal{X}$, and the corresponding label space $\mathcal{Y}$. Each instance $\mathbf{X}_{i}\in\mathbb{R}^{K\times L}$ represents the values of each time series obtained from sensors, where $K$ is the dimensionality of features, and $L$ is the temporal length of the series. Moreover, each instance $\mathbf{X}_{i}$ corresponds to an activity class label $y_{i}\in\{1,2,\cdots,C\}$, indicating the specific activity category performed by the subjects, with $C$ denoting the total number of activity categories. The domain generalization challenge lies in the nonintersection and domain differences between the training and testing sets. Typically, the training set $D^{s}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n^{s}}$ is collected from the labeled source-domain subjects, where $n^{s}$ represents the number of training instances. Importantly, $n^{s}$ is often small in cross-person activity recognition scenarios, presenting the small-scale challenge. On the other hand, the test set $D^{t}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n^{t}}$ consists of $n^{t}$ instances obtained from the unseen target-domain subjects and satisfies the condition $D^{s}\cap D^{t}=\emptyset$. In addition, source and target domains have different joint probability distributions while sharing the identical feature space and class label space, i.e., $P^{s}(\mathbf{X}_{i},y_{i})\neq P^{t}(\mathbf{X}_{i},y_{i})$, and $\mathcal{X}^{s}=\mathcal{X}^{t}$, $\mathcal{Y}^{s}=\mathcal{Y}^{t}$. The primary objective is to leverage the available data in $D^{s}$ to train a TSC model $f:\mathcal{X}\rightarrow\mathcal{Y}$ capable of effectively generalizing to an inaccessible, unseen test domain $D^{t}$, without any prior exposure to target domain data or domain labels during training. This task is inherently more challenging than conventional transfer learning settings (Qin et al., 2023; Qian et al., 2021) due to the disparate distributions across source and target domains, compounded by the small-scale settings of the training data.
@@ 3.2. Diffusion Probabilistic Model
Sampling procedure.   Given a well-trained $p_{\theta}$, the data generation procedure begins with a Gaussian noise $x_{T}\sim\mathcal{N}(0,\textbf{I})$ and proceeds by iteratively denoising $x_{t}$ for $t=T,\cdots,1$ through $p_{\theta}(x_{t-1}|x_{t})$, culminating in the generation of the new data $x_{0}$.
@@ 5.1. Activity Style Condition
Conditional diffusion models are typically guided by label or text prompts that provide task-specific knowledge, such as ”create a [cartoonish] [cat] image” (Min et al., 2023; Zhang et al., 2022b; Wang et al., 2022b). However, the generation of instance-level time-series data introduces distinct challenges. It is difficult to capture the complex patterns solely through label or text prompts due to the inherently high-dimensional and non-stationary nature (Huang et al., 2020b). To address this issue, we propose the development of a style conditioner using a contrastive learning approach (Eldele et al., 2021). This approach has demonstrated robustness in extracting representations from unlabeled time-series data. The transformed data can retain the distinctive characteristics of the original data while preserving the semantic information of the classes. Thus, it is well-suited for extracting robust instance-level representation, termed as “style”, which can serve as conditions to guide diffusion models.
Delving into specifics, the contrastive learning pipeline consists of a feature encoder and Transformer on the available training data. The objective is to maximize the similarity between different contexts of the same sample and minimize the similarity between contexts of different samples. Once the module is trained, we utilize it as style conditioner denoted as $f_{\text{style}}$. When extracting the style from the original data $\mathbf{X}_{i}$, the style conditioner produces a style vector $S_{i}=f_{\text{style}}(\mathbf{X}_{i})\in\mathbb{R}^{H}$, where $H$ denotes the length of the vector. Consequently, each activity style condition can be interpreted as ”a [$y_{i}$] activity performed in [${S}_{i}$] style”, where $y_{i}$ denotes the class of the original data. This approach takes an important step towards the first criteria of domain padding due to the preservation of class semantics. The aggregation of all context vectors from $n^{s}$ training instances constitutes a set $\mathcal{S}=\{{S_{i}}\}_{i=1}^{n^{s}}$, which can be further divided into $C$ class-specific subsets corresponding to $C$ classes. Each subset contains style vectors pertaining to a specific class, expressed as $\mathcal{S}=\{\mathcal{S}^{1}\cup\mathcal{S}^{2}\cup\dots\cup\mathcal{S}^{C}\}$. In Appendix A, we provide the details of the contrastive learning approach (Eldele et al., 2021).
@@ 5.3. Beyond One Activity Style
(7) $\mathcal{P}(\mathcal{S}^{c})=\{\mathcal{D}_{j}|\mathcal{D}_{j}\subseteq\mathcal{S}^{c},\mathcal{D}_{j}\neq\emptyset\}$
@@ 6.1. Experimental Setup
Table 1. Classification accuracy (%) ($\pm$ standard deviation) on three public datasets, where each task only comprises 20% of training data. The best results are marked in bold. ”T0-T4” represent different cross-person activity recognition tasks.
@@ 6.4. Ablation and Sensitivity Analysis
In this section, we perform an ablation study that focuses on the main step of DI2SDiff, i.e., generating diverse time-series data via diffusion for data augmentation. We keep the number of synthetic samples and the training strategy of TSC models the same for all variants. Additionally, we conduct a sensitivity analysis that focuses on its two hyperparameters: $\kappa$, which controls the volume of synthetic data, and $o$, which controls the maximum number of style features in each style combination.
@@ ACKNOWLEDGEMENTS
This work was supported in part by Zhejiang Science and Technology Plan Project under Grant 2023C03183, Key Scientific Research Base for Digital Conservation of Cave Temples (Zhejiang University), China NSFC under Grant 62172284, and Guangdong Basic and Applied Basic Research Foundation under Grant 2022A1515010155.
@@ D.1. Details of training TSC models
Overall. During the training process, as Algorithm 1 shows, the three steps are iteratively repeated in each epoch until termination. In the inference phase, as Algorithm 2 shows, we only use the trained feature extractor $G_{f}$, projection layer $G_{\text{proj}^{\text{c}}}$ and the class classifier $G_{y^{\text{c}}}$ from step (iii). They are stacked together to classify the input time-series samples in the test dataset $D^{t}$.

### v2

@@ Abstract.
Existing domain generalization (DG) methods for cross-person generalization tasks often face challenges in capturing intra- and inter-domain style diversity, resulting in domain gaps with the target domain. In this study, we explore a novel perspective to tackle this problem, a process conceptualized as domain padding. This proposal aims to enrich the domain diversity by synthesizing intra- and inter-domain style data while maintaining robustness to class labels. We instantiate this concept using a conditional diffusion model and introduce a style-fused sampling strategy to enhance data generation diversity. In contrast to traditional condition-guided sampling, our style-fused sampling strategy allows for the flexible use of one or more random styles to guide data synthesis. This feature presents a notable advancement: it allows for the maximum utilization of possible permutations and combinations among existing styles to generate a broad spectrum of new style instances. Empirical evaluations on a broad range of datasets demonstrate that our generated data achieves remarkable diversity within the domain space. Both intra- and inter-domain generated data have proven to be significant and valuable, contributing to varying degrees of performance enhancements. Notably, our approach outperforms state-of-the-art DG methods in all human activity recognition tasks.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: XXXXXXX.XXXXXXXisbn: 979-8-4007-0490-1/24/08ccs: Human-centered computing Ubiquitous computingccs: Computing methodologies Transfer learning
@@
Figure 1. T-SNE visualization of time-series features extracted by various methods across three domains in HAR. Existing representation learning methods result in domain gaps as in both (a) and (b), covering a small portion of target domain (red circles). Standard data augmentation (DA) leads to augmented data (stars), with source domains (orange/blue circles) remaining in close proximity to each other and failing to fill gaps. Our method (d) creates a comprehensive feature space by padding domain gaps via the idea of (e).
@@ 2. Related Work
Human activity recognition (HAR) uses wearable sensors for recognizing activities in healthcare and human-computer interaction (Chen et al., 2021; Xu et al., 2023). The complexity of daily activities, varying among individuals with different personal styles, makes recognition challenging (Qian et al., 2021; Zhang et al., 2022a). With the rise of deep learning, deep neural networks have been increasingly utilized to extract informative features from activity signals (Yang et al., 2015; Jiang and Yin, 2015; Zhang et al., 2022a, 2023a). For example, DeepConvLSTM incorporates convolutional and LSTM units for multimodal wearable sensors (Qiao et al., 2020). MultitaskLSTM extracts features using shared weights, then classifies activities and estimates intensity separately (Barut et al., 2020).
@@ 3.1. Problem Statement
In cross-person activity recognition (Qin et al., 2023), a domain is characterized by a joint probability distribution $P_{X,Y}$ across the product space of time-series instances $\mathcal{X}$, and the corresponding label space $\mathcal{Y}$. Each instance $\mathbf{X}_{i}\in\mathbb{R}^{K\times L}$ represents the values of each time series obtained from sensors, where $K$ is the dimensionality of features, and $L$ is the temporal length of the series. Moreover, each instance $\mathbf{X}_{i}$ corresponds to an activity class label $y_{i}\in\{1,2,\dots,C\}$, indicating the specific activity category performed by the subjects, with $C$ denoting the total number of activity categories. The domain generalization challenge lies in the nonintersection and domain differences between the training and testing sets. Typically, the training set $D^{s}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n^{s}}$ is collected from the labeled source-domain subjects, where $n^{s}$ represents the number of training instances. Importantly, $n^{s}$ is often small in cross-person activity recognition scenarios, presenting the small-scale challenge. On the other hand, the test set $D^{t}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n^{t}}$ consists of $n^{t}$ instances obtained from the unseen target-domain subjects and satisfies the condition $D^{s}\cap D^{t}=\emptyset$. In addition, source and target domains have different joint probability distributions while sharing the identical feature space and class label space, i.e., $P^{s}(\mathbf{X}_{i},y_{i})\neq P^{t}(\mathbf{X}_{i},y_{i})$, and $\mathcal{X}^{s}=\mathcal{X}^{t}$, $\mathcal{Y}^{s}=\mathcal{Y}^{t}$. The primary objective is to leverage the available data in $D^{s}$ to train a TSC model $f:\mathcal{X}\rightarrow\mathcal{Y}$ capable of effectively generalizing to an inaccessible, unseen test domain $D^{t}$, without any prior exposure to target domain data or domain labels during training. This task is inherently more challenging than conventional transfer learning settings (Qin et al., 2023; Qian et al., 2021) due to the disparate distributions across source and target domains, compounded by the small-scale settings of the training data.
@@ 3.2. Diffusion Probabilistic Model
Sampling procedure.   Given a well-trained $p_{\theta}$, the data generation procedure begins with a Gaussian noise $x_{T}\sim\mathcal{N}(0,\textbf{I})$ and proceeds by iteratively denoising $x_{t}$ for $t=T,\dots,1$ through $p_{\theta}(x_{t-1}|x_{t})$, culminating in the generation of the new data $x_{0}$.
@@ 5.1. Activity Style Condition
Conditional diffusion models are typically guided by label or text prompts that provide task-specific knowledge, such as “create a [cartoonish] [cat] image” (Min et al., 2023; Zhang et al., 2022b; Wang et al., 2022b). However, the generation of instance-level time-series data introduces distinct challenges. It is difficult to capture the complex patterns solely through label or text prompts due to the inherently high-dimensional and non-stationary nature (Huang et al., 2020b). To address this issue, we propose the development of a style conditioner using a contrastive learning approach (Eldele et al., 2021). This approach has demonstrated robustness in extracting representations from unlabeled time-series data. The transformed data can retain the distinctive characteristics of the original data while preserving the semantic information of the classes. Thus, it is well-suited for extracting robust instance-level representation, termed as “style”, which can serve as conditions to guide diffusion models.
Delving into specifics, the contrastive learning pipeline consists of a feature encoder and Transformer on the available training data. The objective is to maximize the similarity between different contexts of the same sample and minimize the similarity between contexts of different samples. Once the module is trained, we utilize it as style conditioner denoted as $f_{\text{style}}$. When extracting the style from the original data $\mathbf{X}_{i}$, the style conditioner produces a style vector $S_{i}=f_{\text{style}}(\mathbf{X}_{i})\in\mathbb{R}^{H}$, where $H$ denotes the length of the vector. Consequently, each activity style condition can be interpreted as “a [$y_{i}$] activity performed in [${S}_{i}$] style”, where $y_{i}$ denotes the class of the original data. This approach takes an important step towards the first criteria of domain padding due to the preservation of class semantics. The aggregation of all context vectors from $n^{s}$ training instances constitutes a set $\mathcal{S}=\{{S_{i}}\}_{i=1}^{n^{s}}$, which can be further divided into $C$ class-specific subsets corresponding to $C$ classes. Each subset contains style vectors pertaining to a specific class, expressed as $\mathcal{S}=\{\mathcal{S}^{1}\cup\mathcal{S}^{2}\cup\cdots\cup\mathcal{S}^{C}\}$. In Appendix A, we provide the details of the contrastive learning approach (Eldele et al., 2021).
@@ 5.3. Beyond One Activity Style
(7) $\mathcal{P}(\mathcal{S}^{c})=\{\mathcal{D}_{j}|\mathcal{D}_{j}\subseteq\mathcal{S}^{c},\mathcal{D}_{j}\neq\emptyset\}.$
@@ 6.1. Experimental Setup
Table 1. Classification accuracy (%) ($\pm$ standard deviation) on three public datasets, where each task only comprises 20% of training data. The best results are marked in bold. “T0-T” represent different cross-person activity recognition tasks.
@@ 6.4. Ablation and Sensitivity Analysis
In this section, we perform an ablation study that focuses on the main step of DI2SDiff, i.e., generating diverse time-series data via diffusion for data augmentation. We keep the number of synthetic samples and the training strategy of TSC models the same for all variants. Additionally, we conduct a sensitivity analysis that focuses on its two hyperparameters: $o$, which controls the maximum number of style features in each style combination, and $\kappa$, which controls the volume of synthetic data.
@@ ACKNOWLEDGEMENTS
This work is supported by Zhejiang Science and Technology Plan Project under Grant 2023C03183, Key Scientific Research Base for Digital Conservation of Cave Temples (Zhejiang University), China NSFC under Grant 62172284, and Guangdong Basic and Applied Basic Research Foundation under Grant 2022A1515010155.
@@ D.1. Details of training TSC models
Overall.
During the training process, as Algorithm 1 shows, the three steps are iteratively repeated in each epoch until termination. In the inference phase, as Algorithm 2 shows, we only use the trained feature extractor $G_{f}$, projection layer $G_{\text{proj}^{\text{c}}}$ and the class classifier $G_{y^{\text{c}}}$ from step (iii). They are stacked together to classify the input time-series samples in the test dataset $D^{t}$.